40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2493     Entry time: Sat Jan 9 15:02:01 2010
Author: Alberto 
Type: Update 
Category: ABSL 
Subject: PRC scanning 
I scanned the PRC in the frequency range of 30-60 MHz, untill the PLL lost lock. But everything is working fine.
The PRC remained lock for all time, with SPOB at ~1000.
I'm leaving the lab now, planning to come back tomorrow.
I turned the flipping mirrors down and closed the mechanical shutter of the auxiliary NPRO.
ELOG V3.1.3-